A boy rows his boat 8 km downstream in 1 hour and 4 minutes, and returns in 2 2/7 hour at what rate will he row in still water? what is the rate of the current ?
Answer by stanbon(75887)

A boy rows his boat 8 km downstream in 1 hour and 4 minutes, and returns in 2 2/7 hour at what rate will he row in still water? what is the rate of the current ?
------------------------
Downstream DATA:
distance = 8 km ; time = (64/60) hrs ; rate = 8/(64/60) = 7.5 mph
-----
Upstream DATA:
distance = 8 km ; time = 16/7 hrs ; rate = 8/(16/7) = 3.5 mph
------------
Equations:
b + c = 7.5
b - c = 3.5
----
2b = 11 mph
b = 5.5 mph (speed of the boat in still water)
---
Solve for "c":
b + c = 7.5 mph
5.5 + c = 7.5 mph
c = 2 mph (speed of the current)
=================
Cheers,
Stan H.
===================